OpenStudy (anonymous):

Which set of side lengths will form a right triangle? 6, 8, 11 10, 24, 25 7, 24, 26 8, 15, 17

OpenStudy (anonymous):

\(8^2+15^2=17^2\\289=289\)

OpenStudy (sloppycanada):

\[a^2 + b^2 = c^2\]

OpenStudy (anonymous):

thanks <3

OpenStudy (anonymous):

\(\huge \color{pink}{\heartsuit}\)

OpenStudy (anonymous):

;) <3